A monopolist is faced with two distinct damned function given as Q1=48-025P1----(1), Q2=15-0.04P2----(2).If TC=60+15Q----(3) Question. (1) Demonstrate the profit levels for market with discrimination. (2) Demonstrate the profit levels for market without discrimination.(B) Discuss your observation.


1
Expert's answer
2021-07-19T10:20:45-0400

Solution:

1.). The profit levels for the market with discrimination:

We will have to profit maximize the two distinct markets separately:

First derive the profit that maximizes market 1:

Demand function for market 1: Q1 = 48 - 0.25P1

Derive the inverse demand function:

P1 = 192 – 4Q

Derive the total revenue:

TR = "P\times Q = (192 - 4Q)\times Q = 192Q - 4Q^{2}"

Derive Marginal Revenue (MR):


MR = "\frac{\partial TR} {\partial Q} = 192 - 8Q"

Derive Marginal Cost (MC):

TC = 60 + 15Q


MC = "\frac{\partial TC} {\partial Q} = 15"

Set MR = MC to maximize profits:

MR = MC

192 – 8Q = 15

192 – 15 = 8Q

177 = 8Q

Q = 22.125 or 22 units

Substitute in the demand function to derive price:

P1 = 192 – 4Q

P1 = 192 – 4(22) = 192 – 88 = 104

Profit level of market 1:

Profit = TR – TC

Profit = (104"\times"22) – (60 + 15(22))

Profit = 2,288 – 390 = 1,898

Profit level of market 1 = 1,898


Profit maximization for market 2:

Demand function for market 2: Q2 = 15 - 0.04P2

Derive the inverse demand function:

P2 = 375 – 25Q

Derive the total revenue:

TR = P"\times"Q = (375 – 25Q) "\times" Q = 375Q – 25Q2

Derive Marginal Revenue (MR):

MR = "\frac{\partial TR} {\partial Q} = 375 - 50Q"

Derive Marginal Cost (MC):

TC = 60 + 15Q

MC = "\frac{\partial TC} {\partial Q} = 15"

Set MR = MC to maximize profits:

MR = MC

375 – 50Q = 15

375 – 15 = 50Q

360 = 50Q

Q = 7.2 or 7 units

Substitute in the demand function to derive price:

P2 = 375 – 25Q

P2 = 375 – 4(7) = 375 – 28 = 347

Profit level of market 2:

Profit = TR – TC

Profit = (347"\times"7) – (60 + 15(7))

Profit = 2,429 – 165 = 2,264

Profit level of market 2 = 2,264


2.). Demonstrate the profit levels for the market without discrimination:

We have to derive one single price for both markets; therefore, we combine or sum up the demand functions of both markets:

Q1 = 48 - 0.25P1

Q2 = 15 - 0.04P2

QT = 63 – 0.29P

Derive the inverse demand function of the total demand function:

P = 217 – 3Q

Derive TR:

TR = P"\times"Q = (217 – 3Q) "\times" Q = 217Q – 3Q2

Derive Marginal Revenue (MR):

MR = "\frac{\partial TR} {\partial Q} = 217 - 6Q"

Derive Marginal Cost (MC):

TC = 60 + 15Q

MC = "\frac{\partial TC} {\partial Q} = 15"

Set MR = MC to maximize profits:

MR = MC

217 – 6Q = 15

217 – 15 = 6Q

202 = 6Q

Q = 33.7 or 34 units

Substitute in the demand function to derive price:

P = 217 – 3Q

P = 217 – 3(34)

P = 217 – 102 = 115

Combined market single price = 115

Profit level of combined market:

Profit = TR – TC

Profit = (115"\times"34) – (60 + 15(34))

Profit = 3,910 – 570 = 3,340

Profit level of the combined market without price discrimination = 3,340

 

B.). From the above calculations, it is evident that the market generates more profits when there is no price discrimination compared to when there is price discrimination.